40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Aug 14 10:10:56 2020, Jon, Configuration, VAC, Vacuum repairs today 
    Reply  Sat Aug 15 02:02:13 2020, Jon, Configuration, VAC, Vacuum repairs today 
       Reply  Sat Aug 15 15:12:22 2020, Jon, Configuration, VAC, Overhaul of small turbo pump interlocks small_tp_signal_routing.pngsmall_tp_interlock_signal.pngsmall_tp_signal_routing.pngmedm_screen.png
Message ID: 15527     Entry time: Sat Aug 15 02:02:13 2020     In reply to: 15526     Reply to this: 15528
Author: Jon 
Type: Configuration 
Category: VAC 
Subject: Vacuum repairs today 

Vacuum work is completed. The TP2 and TP3 interlocks have been overhauled as proposed in ELOG 15499 and seem to be performing reliably. We're now back in the nominal system state, with TP2 again backing for TP1 and TP3 pumping the annuli. I'll post the full implementation details in the morning.

I did not get to setting up the new UPS units. That will have to be scheduled for another day.

Quote:

The vac system is going down now for planned repairs [ELOG 15499]. It will likely take most of the day. Will advise when it's back up.

ELOG V3.1.3-